A sticker warehouse sells an average of 6 rolls of stickers per customer at $4 per roll. Statistics show that for every $0.25 decrease in price, customers will buy an additional roll. According to this model, at what sticker price will the revenue from stickers be $28.
Answer by Theo(13342)

price * number of items = revenue.
when the price is 4.00, the average number of rolls per customer is 6.
with every .25 decrease in price, the average number rolls per customer goes up 1.
the euation is y = (4 - .25x) * (6 + x).
when x = 0, the equation becomes y = 4 * 6 = 24
when x = 1, the equation becomes y = 3.75 * 7 = 23.25
when x = 2, the equation becomes y = 3.5 * 8 = 28

the revenue will be 28 when the sticker price is 3.5.

here's a graph of the equation.



the graph shows that, when x = 2.
this gets a price per ticket of 3.5 with an average number of 8 tickets per customer.

the maximum revenue occurs when x = 5.
this gets a price per ticket of 2.75 with an average number of 11 tickets per customer.